OpenStudy (anonymous):

simplify radical 1/6

OpenStudy (anonymous):

You need to convert it into a form that has a rationalized denominator. Simply\[\sqrt{\frac{1}{6}}=\frac{1}{\sqrt6}=\frac{1}{\sqrt6}\frac{\sqrt6}{\sqrt6}=\frac{\sqrt6}{6}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

could you help me with some more problems?

OpenStudy (anonymous):

What do you have?

OpenStudy (anonymous):

uhm simplyify the square root of 8x^4y^4

OpenStudy (anonymous):

Factor the argument of the starting expression antd work on them separately. So it goes\[\sqrt{8x^4y^4}=\sqrt8 \sqrt{x^4}\sqrt{y^4}=2\sqrt2x^2y^2\]

OpenStudy (anonymous):

solve 6x^2+x-2=0

OpenStudy (anonymous):

Factor and use the zero product property.